OpenStudy (anonymous):

i need help plss. im not sure how to find the average rate. For f(x) = 0.01(2)x, find the average rate of change from x = 2 to x = 10. 1.275 8 10.2 10.24

jimthompson5910 (jim_thompson5910):

the function is this? \[\LARGE f(x) = 0.01(2)^x\]

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

are you able to find the value of f(2) and f(10) ?

jimthompson5910 (jim_thompson5910):

let me know what you get for those values or let me know if you aren't sure how to find the function values

OpenStudy (anonymous):

im not even sure how to find the function values.. i get all mixed up when it comes to trying to find the answers to these type of questions.

jimthompson5910 (jim_thompson5910):

what you do is replace every x value with the input in question so say you wanted to find f(2) replace every x with 2 \[\LARGE f(\color{red}{x}) = 0.01(2)^\color{red}{x}\] \[\LARGE f(\color{red}{2}) = 0.01(2)^\color{red}{2}\] \[\LARGE f(2) = ???\]

jimthompson5910 (jim_thompson5910):

at this point, you use a calculator to compute `0.01*(2)^2` you can use this calculator if you don't have one http://web2.0calc.com/

OpenStudy (anonymous):

i got 0.04 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

me too

jimthompson5910 (jim_thompson5910):

so f(2) = 0.04

jimthompson5910 (jim_thompson5910):

what is the value of f(10) ?

OpenStudy (anonymous):

100,000,000 ? ._.

jimthompson5910 (jim_thompson5910):

way too big

jimthompson5910 (jim_thompson5910):

did you type in `0.01*(2)^10` ?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

which calculator are you using?

OpenStudy (anonymous):

i got 10.24 this time.. i made a mistake

jimthompson5910 (jim_thompson5910):

that's ok

jimthompson5910 (jim_thompson5910):

now you compute \[\large \frac{f(b)-f(a)}{b-a} = \frac{f(10)-f(2)}{10-2} = \frac{10.24 - 0.04}{10-2} = ??\]

OpenStudy (anonymous):

it equals 1.275

jimthompson5910 (jim_thompson5910):

yep, that's your avg rate of change

OpenStudy (anonymous):

thank you so much for your help! :)

jimthompson5910 (jim_thompson5910):

you're welcome